006 (part 1 of 2) 10.0 points A ladder rests against a vertical l There is no friction between the wall and t adder The coefficient of static friction between the ladder and the ground is0.581 ue 18-19 marder (OlsonM302K1819 2 6 A2, B, CI 7. Al. B C2 8. A2, B1, C3 9. Ai, B2, C3 10. B C3 007 (part 2 of 2) 10.0 points Determine the smallest angle θ for which the ladder remains stationary. Answer in units of

ew -Quest Learning & Assessment Consider the following expressions: where the ground wall, f: force of friction between tbe ladder and F: normal force one ladder due to the 0: angle between the ladder and the ground N: normal force on the der due to the ground W: weight of the ladder, and : length of the ladder Identify the s o qatins which is cor- rect 1. A2, B2. C1 2. A2, B C2 3. A B, CI 4. A B2, CI 5. A B2, C2

the answer to 6 is 9 I just do not know 7
Answer #1

using equilibrium of force along the horizontal direction

f = F                                                eq-1

using equilibrium of force in vertical direction

N = w                                                eq-2

static frictional force is given as

f = \mu _{s} N

using eq-2

f = \mu _{s} w

using eq-1

\mu _{s} w = F                                   eq-3

Using equilibrium of torque about "A"

F Sin\theta (AC) = (w Cos\theta) (AB)

given that : AB = L/2   and AC = L

using eq-3

(\mu _{s} w ) Sin\theta (L) = (w Cos\theta) (L/2)

(\mu _{s} ) Sin\theta = (Cos\theta) (1/2)

Sin\theta /Cos\theta = (0.5) /\mu _{s}

tan\theta = 0.5/0.581

\theta = 40.7 deg